### v2

# Scenario Analysis with Multivariate Bayesian Machine Learning Models

Scenario Analysis with Multivariate Bayesian Machine Learning Models
Michael PFARRHOFER
WU Vienna
Anna STELZER
Oesterreichische Nationalbank
We present an econometric framework that adapts tools for scenario analysis, such as variants of conditional forecasts and impulse response functions, for use with dynamic nonparametric multivariate models. We demonstrate the utility of our approach with simulated data and three real-world applications: (1) scenario-based conditional forecasts aligned with Federal Reserve stress test assumptions, measuring (2) macroeconomic risk under varying financial conditions, and (3) asymmetric effects of US-based financial shocks and their international spillovers. Our results indicate the importance of nonlinearities and asymmetries in dynamic relationships between macroeconomic and financial variables.
JEL: C11, C32, C53, C54
Keywords: conditional forecast, generalized impulse response function, Bayesian additive regression trees, nonlinearities, structural inference
Contact: michael.pfarrhofer@wu.ac.at, Department of Economics, WU Vienna University of Economics and Business. We would like to thank Niko Hauzenberger for useful comments. Disclaimer: The views expressed in this paper do not necessarily reflect those of the Oesterreichische Nationalbank or the Eurosystem.
1. Introduction
In this paper, we discuss how to conduct scenario analysis with multivariate time series models in a macroeconomic context when the functional form of the conditional mean is nonlinear and/or unknown. This situation can arise when relying on “traditional” nonlinear frameworks (e.g., in variants of threshold, regime-switching, or time-varying parameter models, see Fischer et al. (2023) for a recent example), but is virtually always the case in recently developed models that introduce Bayesian machine learning (ML) techniques to multivariate macroeconometric modeling (see, e.g., Huber and Rossini, 2022; Clark et al., 2023; Huber et al., 2023; Hauzenberger et al., 2024c). Our framework combines elements of structural vector autoregressions (SVARs) with nonlinear reduced form models, which is why we sometimes refer to aspects of our approach as “semi-structural.”
We use the term scenario analysis rather broadly to refer to different counterfactual experiments, including a version of conditional forecasts (CFs) and variants of nonlinear impulse response functions (IRFs). While some aspects and issues in the Bayesian ML context have been discussed in isolation in the aforementioned papers, there is no unified explicit treatment or comprehensive framework available yet. Inspired by the related work of Crump et al. (2025), which focuses on a linear Bayesian vector autoregression (VAR), this is the gap in the literature that we aim to fill with our paper.
CFs — simulating the future path of a subset of variables conditional on predefined scenarios encoded in constraints placed on other observed variables — have been used by academics and practitioners since the 1980s in a linear VAR context (see, e.g., Doan et al., 1984), with subsequent refinements and improvements in computational efficiency (see, e.g., Waggoner and Zha, 1999; Jarociński, 2010; Bańbura et al., 2015). A recent and highly efficient computational approach is developed in Chan et al. (2025). Their framework exploits the properties of conditionally Gaussian errors in conjunction with the common assumption of a known linear conditional mean relationship.
Breaking the assumption of linearity complicates matters. In particular, there is no general causal representation of nonlinear time series as functions of structural shocks (see, e.g., Potter, 2000, for a discussion), and the presence of unknown nonlinearities makes it difficult to derive multi-step ahead predictive distributions. In many cases it is outright impossible to obtain closed-form solutions. In this case, one can resort to recursive predictive simulations. Indeed, this paper discusses how common approaches to (conditional) forecasting can be adapted to the context of nonparametric multivariate models for various types of scenario analyses based on Monte Carlo methods. The proposed approach is quite general in the sense that it can be used in conjunction with most popular frameworks that introduce nonlinearities in the conditional mean of multivariate models. The key assumption is that the errors are conditionally multivariate Gaussian.
Since the IRF, another key macroeconomic estimand, can be defined as the difference between forecasts conditioned on the values of different shocks (for seminal work, see Gallant et al., 1993; Koop et al., 1996), our proposed framework is also useful for nonlinear semi-structural scenario analysis. For excellent discussions in the context of linear SVARs and partially nonlinear structural models, see Antolin-Diaz et al. (2021) and Gonçalves et al. (2021). Specifically, we revisit the issue of how to obtain dynamic causal effects, in the form of generalized IRFs (GIRFs), in models with nonlinearities of unknown form. Compared with IRFs from standard linear VARs, which are symmetric, shape invariant and history independent, our GIRFs do not feature these potentially restrictive properties. We examine both unrestricted GIRFs to shocks identified with approaches typically used in the linear SVAR literature, and restricted GIRFs, which can be used to investigate and quantify the contributions of specific sets of transmission channels in the propagation of shocks. The latter are obtained by partially matching the moments of the unconditional and conditional predictive distributions used to compute the GIRFs.
We mentioned above that the proposed approach is quite generally applicable in multivariate dynamic models. This means that there is a large number of potential candidates for the respective conditional mean functions that can be assumed. These choices include (but are not limited to) regression tree-based implementations or Gaussian process priors (see Marcellino and Pfarrhofer, 2024, for a recent review). For our empirical work, we use Bayesian Additive Regression Trees (BART, Chipman et al., 2010) as a specific nonparametric implementation to infer functional relationships in a multivariate time series model. We pick this sum-of-trees model because tree-based approaches have proven particularly capable of producing accurate forecasts when used with time series data for the US economy (see, e.g., Medeiros et al., 2021; Goulet Coulombe et al., 2022; Clark et al., 2023; Goulet Coulombe, 2024), with datasets typically structured similarly to the one we use in this paper.
Our approach to scenario analysis is developed under the assumption of a multivariate Gaussian reduced form error term with a general time-varying covariance matrix. From a practical and implementation perspective, we rely on a somewhat simplified version to capture time variation of the respective volatilities. The framework allows for flexible equation-by-equation estimation of the multivariate system and offers significant improvements in computational efficiency also in the context of CFs and GIRFs. In our empirical work, we capture heteroskedastic features of the data with a specification related to common volatility approaches (see Carriero et al., 2016), which reflects recent tools used to address outliers during the Covid-19 pandemic (see, e.g., Carriero et al., 2024). It is worth noting that all methods also work with more sophisticated volatility models (see, e.g., Chan, 2023), albeit at the cost of an increased computational burden.
After briefly investigating the performance of our approach with synthetic data, we apply our framework in several related yet distinct cases. Our dataset comprises about $25$ quarterly macroeconomic and financial variables for the US economy ranging from the mid-$1970$s to the last quarter of $2023$. In one of our explorations, we also add some international variables and include data from the euro area (EA), Japan, and the United Kingdom (UK) in our model. The applied work assesses and illustrates the role of nonlinearities when interest centers on CFs, and we also explore asymmetries in the propagation of shocks of different types, signs and magnitudes. Specifically, we provide three empirical applications. First, inspired by Chan et al. (2025), we use a subset of the assumptions underlying the annual stress test conducted by the Federal Reserve System and compute CFs using soft constraints for different scenarios, comparing predictive densities from linear and nonlinear models. Second, reflecting the growth-at-risk literature (see Adrian et al., 2019), we study the unorthogonalized counterfactual implications of varying financial conditions (imposed via a hard constraint) on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock (as in, e.g., Barnichon et al., 2022), and compute GIRFs to shocks of different signs and magnitudes that are allowed to propagate internationally. We then use a restricted GIRF approach to gauge the role of spillovers and spillbacks.
The rest of this paper is structured as follows. Section 2 discusses challenges and solutions for how to obtain predictive inference in the presence of nonlinearities of unknown form. We discuss how to impose constraints on forecasts, and how these constraints may be used to construct scenarios through GIRFs. Section 3 discusses an econometric implementation using BART. Section 4 contains three empirical illustrations. The last section concludes.
2. Nonlinearity and Predictive distributions
Let $\bm{y}_{t}=(y_{1t},\ldots,y_{nt})^{\prime}$ collect $n$ variables for $t=1,\ldots,T,$ and $\bm{x}_{t}=(\bm{y}_{t-1}^{\prime},\ldots,\bm{y}_{t-p}^{\prime})^{\prime}$ is a $k=np$ vector of lags. Interest centers on multivariate models of the form:
$\bm{y}_{t}=\bm{F}(\bm{x}_{t})+\bm{\epsilon}_{t},\quad\bm{\epsilon}_{t}\sim\mathcal{N}(\bm{0}_{n},\bm{\Sigma}_{t}),$ (1)
where $\bm{F}(\bm{x}_{t})=(f_{1}(\bm{x}_{t}),\ldots,f_{n}(\bm{x}_{t}))^{\prime}$ is an $n$-vector of conditional mean functions $f_{i}(\bm{x}_{t}):\mathbb{R}^{k}\rightarrow\mathbb{R}$ for $i=1,\ldots,n,$ such that $\bm{F}(\bm{x}_{t}):\mathbb{R}^{k}\rightarrow\mathbb{R}^{n}$. We assume iid reduced form Gaussian errors $\bm{\epsilon}_{t}$, with $n\times n$ time-varying covariance matrix $\bm{\Sigma}_{t}$. [^1] For the most part we rely on reduced form modeling and introduce additional structure only selectively. A feasible alternative that comes with advantages and disadvantages (see Section 3), is to parameterize Equation (1) directly as $\bm{y}_{t}=\bm{F}(\bm{x}_{t})+\bm{B}_{0}^{-1}\bm{u}_{t},~{}\bm{u}_{t}\sim\mathcal{N}(\bm{0}_{n},\bm{H}_{t})$, where $\bm{B}_{0}$ is nonsingular (and possibly lower triangular) and $\bm{H}_{t}$ is a diagonal matrix, see also Arias et al. (2023); Chan et al. (2024). One may assume a specific functional form for the $f_{i}(\bm{x}_{t})$’s or treat them as unknown and estimate them. The methods we propose in this paper are designed specifically for the latter case and we discuss one (of several) possible model implementations in Section 3.2.
Many of the tools for the various kinds of scenario analysis we discuss in this paper rely on computing functions of the (conditional) moments of $\bm{y}_{t+h}$ based on Equation (1). For expository purposes and to keep the notation simple, consider the single-lag case $p=1$ and $\bm{\Sigma}_{t}=\bm{\Sigma}$, without loss of generality. We may write:
$\bm{y}_{t+h}=\tilde{\bm{F}}_{(h)}(\bm{y}_{t},\bm{\epsilon}_{t+1},\bm{\epsilon}_{t+2},\ldots,\bm{\epsilon}_{t+h})=\bm{F}(\tilde{\bm{F}}_{(h-1)}(\bm{y}_{t},\bm{\epsilon}_{t+1},\ldots,\bm{\epsilon}_{t+(h-1)}))+\bm{\epsilon}_{t+h},$ (2)
with $\bm{y}_{t+1}=\tilde{\bm{F}}_{(1)}(\bm{y}_{t},\bm{\epsilon}_{t+1})=\bm{F}(\bm{y}_{t})+\bm{\epsilon}_{t+1}$, and $\tilde{\bm{F}}_{(h)}(\bullet)$ denotes the $h$-step composition of $\bm{F}(\bullet)$ which is defined recursively. This expression is related to the Wold decomposition, and expresses $\bm{y}_{t+h}$ as a function of the initial condition $\bm{y}_{t}$ and a sequence of white noise shocks, $\{\bm{\epsilon}_{t+1},\ldots,\bm{\epsilon}_{t+h}\}$. [^2] Under some assumptions about the behavior of $\tilde{\bm{F}}_{(h)}(\bullet)$, it is noteworthy that expansions such as the Volterra series could be used as an approximation and a close equivalent to the Wold representation for nonlinear time series, see, e.g., Potter (2000, Section 4) and Jordà (2005, Section II). Our approach does not rely on approximations but resorts to simulation-based methods.
[…]
2.1. Predictive Simulation
Define a vector $\bm{\Xi}$ that contains all coefficients and latent variables necessary to parameterize Equation (1). At time $\tau$, the one-step-ahead predictive distribution is:
[…]
where $\mathcal{I}$ denotes the respective information set used to infer $\bm{\Xi}$. For typically out-of-sample forecast exercises, $\mathcal{I}$ is given by $\{\bm{y}_{t}\}_{t=1}^{\tau}$, and we are interested in predicting $\bm{y}_{\tau+h}\>|\>\{\bm{y}_{t}\}_{t=1}^{\tau}$ for $h=1,2,\ldots,$ steps ahead. In other cases we may also want to condition on the full information set, i.e., $\mathcal{I}$ is given by $\{\bm{y}_{t}\}_{t=1}^{T}$, to compute scenarios in-sample for $\tau\in\{1,2,\ldots,T\}$ using the distribution of $\bm{y}_{\tau+h}\>|\>\{\bm{y}_{t}\}_{t=1}^{T}$ conditional on parameters informed by the full information set. In either case, $p(\bm{y}_{\tau+1}\>|\>\mathcal{I})$ generally does not take a well-known form, and neither does the distribution of higher-order forecasts, for $h\geq 2$.
However, we may still explore and obtain random samples from them via predictive simulation. This involves exploiting the fact that even though $p(\bm{y}_{\tau+1}\>|\>\mathcal{I})$ is unknown, $p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\bm{\Xi})$ takes a conditionally Gaussian form under the model in Equation (1). [^3] In fact, Gaussianity is not a necessary requirement for exploring the unconditional multi-step predictive distribution here (it is, however, for the CFs in Section 2.2). In the context of Equation (4), it is sufficient that the one-step-ahead predictive distribution is either of known form conditionally, or that a sample from it can be generated. Let $m$ denote the current iteration of the MCMC algorithm and $x^{(m)}$ indicates the $m$th draw of a random variable. The one-step-ahead predictive distribution is:
$p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\bm{\Xi}^{(m)})=\mathcal{N}(\bm{F}^{(m)}(\bm{x}_{\tau+1}),\bm{\Sigma}_{\tau+1}^{(m)}),$ (4)
where $\bm{x}_{\tau+1}=(\bm{y}_{\tau}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$. For $h\geq 2$ we iterate forward, conditioning recursively on the draws for preceding horizons, by setting the predictors to $\bm{x}_{\tau+h}^{(m)}=(\bm{y}_{\tau+h-1}^{(m)}{{}^{\prime}},\bm{y}_{\tau+h-2}^{(m)}{{}^{\prime}},\ldots)^{\prime}$, and obtain:
$p(\bm{y}_{\tau+h}\>|\>\mathcal{I},\bm{y}_{\tau+1:\tau+h-1}^{(m)},\bm{\Xi}^{(m)})=\mathcal{N}(\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m)}),\bm{\Sigma}_{\tau+h}^{(m)}),$ (5)
where $\bm{y}_{\tau+1:\tau+h-1}^{(m)}$ denotes the path of the variables from $\tau+1$ to $\tau+h-1$ and $\bm{y}_{\tau+1:\tau+h}=(\bm{y}_{\tau+1}^{\prime},\ldots,\bm{y}_{\tau+h}^{\prime})^{\prime}$. This is equivalent to the recursive composition discussed in the context Equation (2), and exploits the fact that:
$p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I})=\int p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\bm{\Xi})\prod_{j=2}^{h}p(\bm{y}_{\tau+j}\>|\>\bm{y}_{\tau+1:\tau+j-1},\mathcal{I},\bm{\Xi})p(\bm{\Xi}\>|\>\mathcal{I})d\bm{\Xi},$ (6)
that is, the joint distribution of forecasts can be decomposed into the product of the conditional one-step ahead predictive densities. Simulating the process forward, sampling from the distribution in Equation (5) across horizons $h=1,2,\ldots,$ in each sweep of an MCMC algorithm, delivers draws from $p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I})$ via a Monte Carlo approach. It is worth noting here that we may compute posteriors of any functions that take the forecast path as an input. We provide additional discussions in this context in Section 2.3.
2.2. Conditional Forecasts
Suppose we want to impose restrictions on a predefined path of one or more variables. Formally, this implies an additional conditioning argument for the predictive distribution, see also Crump et al. (2025) for related discussions. In this context, we denote by $\mathcal{C}_{h}$ a set that defines the desired restrictions at horizon $h=1,2,\ldots,$ i.e., the unconditional forecast results when $\mathcal{C}_{h}=\varnothing$ for all $h$, and $\mathcal{C}_{1:h}=\{\mathcal{C}_{1},\ldots,\mathcal{C}_{h}\}$. In our version, we condition on observables, and interest centers on the conditional distribution $p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1:h})$, which in line with Equation (6) can be written as:
[…]
$p(\bm{y}_{\tau+1:\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1:h},\bm{\Xi})$ $=p(\bm{y}_{\tau+1}\>|\>\mathcal{I},\mathcal{C}_{1},\bm{\Xi})\prod_{j=2}^{h}p(\bm{y}_{\tau+j}\>|\>\bm{y}_{\tau+1:\tau+j-1},\mathcal{I},\mathcal{C}_{1:j},\bm{\Xi}).$
[…]
We note that this differs from the “traditional” implementation of CFs, that impose the conditions as $p(\bm{y}_{t+j}\>|\>\mathcal{C}_{1:h},\bullet)$ instead of $p(\bm{y}_{t+j}\>|\>\mathcal{C}_{1:j},\bullet)$ for $j=1,\ldots,h$. Put simply, our approach “filters” forward in a loose sense of the word by conditioning on the history of restrictions, thereby resulting in a future path of the unrestricted variables that is consistent with the imposed restrictions over the full set of horizons. By contrast, versions in the spirit of Waggoner and Zha (1999) condition on future restrictions as well as past (and current) ones at each horizon, either by drawing the entire paths of the shocks or by smoothing via backwards recursions (see Bańbura et al., 2015, for an excellent discussion).
In static multivariate problems these two approaches coincide (i.e., when the joint predictive distribution across horizons factors into the product of the marginals), while, intuitively, they will potentially lead to increasingly different results as the persistence of the underlying dynamic processes increases. We explore this apparent limitation and the practical usefulness of our approach below in an exercise using artificial data. For dynamic systems featuring a small to moderate amount of persistence (e.g., as in typical applications with monthly/quarterly macroeconomic/financial data that are transformed towards approximate stationarity), the two versions’ outputs are usually close.
Constraints. Define the selection matrix $\bm{R}_{h}$ of size $r_{h}\times n$, where $r_{h}$ is the number of restrictions which may vary across horizons. This matrix serves to select the respective restricted variables. There are two main cases of importance: hard and soft constraints. In the former case, some variables exactly follow a predetermined path; in the latter, they lie within a predetermined interval. As this interval narrows, the soft-constrained case approaches the hard-constrained one. We store these restrictions in the vector $\bm{r}_{h}$ of size $r_{h}\times 1$, whereas $\underline{\bm{r}_{h}}$ and $\overline{\bm{r}_{h}}$, both of size $r_{h}\times 1$, denote the lower and upper bounds of an interval. For hard constraints, we have an equality restriction (see Equation 7), while the soft constraints require an inequality restriction (see Equation 8):
$\mathcal{C}_{h}$ =$=\{\bm{R}_{h}\bm{y}_{\tau+h}=\bm{r}_{h}\},$ (7)
$\mathcal{C}_{h}$ =$=\{\underline{\bm{r}_{h}}\leq\bm{R}_{h}\bm{y}_{\tau+h}\leq\overline{\bm{r}_{h}}\}.$ (8)
To simplify notation, we omit the $m$-superscript that labels MCMC draws in what follows, but stress that the following computations are carried out in each sweep of our sampling algorithm, thereby marginalizing over the parameters. In addition, define the matrix $\bm{U}_{h}$ of dimension $(n-r_{h})\times n$, which mirrors the selection matrix $\bm{R}_{h}$ and is used to select all unrestricted variables. Define the restricted, $\bm{y}_{\tau+h}^{(R)}=\bm{R}_{h}\bm{y}_{\tau+h}$, and unrestricted, $\bm{y}_{\tau+h}^{(U)}=\bm{U}_{h}\bm{y}_{\tau+h}$, subvectors of $\bm{y}_{\tau+h}$. Similar to Chan et al. (2023), using Equation (5), we may then write:
$\bm{y}_{\tau+h}=$ $\bm{R}_{h}^{\prime}\bm{y}_{\tau+h}^{(R)}+\bm{U}_{h}^{\prime}\bm{y}_{\tau+h}^{(U)},$
$\bm{R}_{h}^{\prime}\bm{y}_{\tau+h}^{(R)}+\bm{U}_{h}^{\prime}\bm{y}_{\tau+h}^{(U)}\sim\mathcal{N}(\bm{F}(\bm{x}_{\tau+h}),\bm{\Sigma}_{\tau+h}).$
This representation can be used to obtain the conditional distribution $p(\bm{y}_{\tau+h}^{(U)}\>|\>\bm{y}_{\tau+h}^{(R)}=\bm{r}_{h},\bullet)$, which is proportional to:
$\exp\left\{-\frac{1}{2}(\bm{y}_{\tau+h}^{(U)}{{}^{\prime}}\bm{U}_{h}\bm{\Sigma}_{\tau+h}^{-1}\bm{U}_{h}^{\prime}\bm{y}_{\tau+h}^{(U)}-2\bm{y}_{\tau+h}^{(U)}{{}^{\prime}}\bm{U}_{h}\bm{\Sigma}_{\tau+h}^{-1}(\bm{F}(\bm{x}_{\tau+h})-\bm{R}_{h}^{\prime}\bm{y}_{\tau+h}^{(R)}))\right\}.$
This is the kernel of a Gaussian distribution, so under the hard restriction in Equation (7), we obtain:
$\bm{y}_{\tau+h}^{(U)}\>|\>\bm{y}_{\tau+h}^{(R)}=\bm{r}_{h},\bullet\sim\mathcal{N}(\bm{m}^{(U)}_{\tau+h},\bm{S}^{(U)}_{\tau+h}),$ (9)
with moments $\bm{S}^{(U)}_{\tau+h}=(\bm{U}_{h}\bm{\Sigma}_{\tau+h}^{-1}\bm{U}_{h}^{\prime})^{-1}$ and $\bm{m}^{(U)}_{\tau+h}=\bm{S}^{(U)}_{\tau+h}(\bm{U}_{h}\bm{\Sigma}_{\tau+h}^{-1}(\bm{F}(\bm{x}_{\tau+h})-\bm{R}_{h}^{\prime}\bm{y}_{\tau+h}^{(R)}))$.
The inequality restrictions of Equation (8) imply that the CF of interest, $p(\bm{y}_{\tau+h}\>|\>\underline{\bm{r}_{h}}\leq\bm{y}_{\tau+h}^{(R)}\leq\overline{\bm{r}_{h}},\bullet)$, follows a truncated multivariate Gaussian. Typically only a subset of variables is restricted (i.e., $r_{h}<n$). Chan et al. (2025) suggest blocked updates to unlock computational advantages — one may decompose the joint distribution of the restricted and unrestricted variables into two parts, and first sample from the $r_{h}$-dimensional (rather than $n$-dimensional) truncated multivariate Gaussian:
$\bm{y}_{\tau+h}^{(R)}\>|\>\underline{\bm{r}_{h}}\leq\bm{y}_{\tau+h}^{(R)}\leq\overline{\bm{r}_{h}},\bullet\sim\mathcal{N}(\bm{m}^{(R)}_{\tau+h},\bm{S}^{(R)}_{\tau+h})\cdot\mathbb{I}(\underline{\bm{r}_{h}}\leq\bm{y}_{\tau+h}^{(R)}\leq\overline{\bm{r}_{h}}),$
with moments $\bm{S}^{(R)}_{\tau+h}=\bm{R}_{h}\bm{\Sigma}_{\tau+h}\bm{R}_{h}^{\prime}$ and $\bm{m}^{(R)}_{\tau+h}=\bm{R}_{h}\bm{F}(\bm{x}_{\tau+h})$. An iid draw from this distribution, $\bm{y}_{\tau+h}^{(R,m)}$, can be obtained via the method of Botev (2017). On may then set $\bm{r}_{h}=\bm{y}_{\tau+h}^{(R,m)}$, and use Equation (9) to draw the unconstrained subvector of forecasts.
(a) hard constraints, moderate persistence
(b) soft constraint, moderate persistence
(c) hard constraints, high persistence
(d) soft constraint, high persistence
Figure 1: Conditional forecast (CF) under moderate and high levels of persistence of the underlying data generating process process (DGP).
Notes: Variables 1, 2, 3. Lines refer to the percentiles 16/50/84 of the predictive distributions (solid grey: recursive simulation; dashed navy blue: precision sampling). The DGP is a VAR with $n=6$ variables and $p=4$ lags. The scenario hard constraints conditions on the realizations of three variables and leaves the others unrestricted; soft contraint restricts one of the variables to a predefined interval.
Simulation. To assess the performance and implications of our recursive approach relative to traditional implementations of CFs, we provide a comparison with the precision sampler of Chan et al. (2025). Assuming a linear conditional mean function allows to contrast both algorithms one-to-one (we cannot use the precision sampler for nonlinear implementations). For this purpose, we simulate artificial data from a linear VAR process with $n=6$, $p=4$ and a maximum horizon $h=20$ for different levels of persistence (measured through the spectral radius $\rho(\tilde{\bm{A}})$, i.e., the maximum absolute eigenvalues of the companion matrix of the VAR, $\tilde{\bm{A}}$). We label $\rho(\tilde{\bm{A}})\approx 0.6$ as moderate persistence, and $\rho(\tilde{\bm{A}})\approx 0.95$ as high persistence. The scenario hard constraints treats the first three variables as missing and conditions on the realizations of the remaining variables; soft contraint restricts one of the variables to the interval $(\mu_{y}-0.1,\mu_{y}+0.1)$ where $\mu_{y}=1/h\sum_{t=T-h+1}^{T}y_{i,t}$.
The resulting CFs are shown in Figure 1. The solid grey lines mark the 68 percent credible set and the median of the predictive distribution when using our proposed recursive approach; the dashed navy blue lines show the same percentiles when using the precision sampler. Considering panels (a) and (b), when facing a moderate amount of persistence, the two approaches yield virtually identical estimates. Some deviations are visible in panels (c) and (d), especially for the left-most panel under the hard constraints and for the first two variables under the soft restriction. In the majority of our experiments, the differences were practically negligible.
2.3. Generalized Impulse Response Functions
The IRF is a ubiquitous object of interest in macroeconomics and widely used for both academic and policy analysis. For the purposes of our paper, we follow the recent literature (see, e.g., Jordà and Taylor, 2024) and define IRFs generically as:
$\text{IRF}_{\tau,h}^{(d,d_{0})}\equiv\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},u_{j\tau}=d_{0}+d)-\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},u_{j\tau}=d_{0})=\bm{\delta}_{\tau,h}^{(d,d_{0})},$ (10)
the difference between two distinct forecasts. As their name suggests, IRFs trace the dynamic evolution of one or more variables in response to a shock impact (the “impulse”) of magnitude $d$ at time $\tau$ across horizons $h$, where the (structural) shock is encoded in $u_{j\tau}$. Note choices about the baseline level of the shock, $d_{0}$, may matter when modeling nonlinear shock impacts. We abstract from this dimension of asymmetry and typically set $d_{0}=0$. Thus we do not indicate the implicit conditioning on the baseline in our notation for any related variables or parameters henceforth. For instance, we simply refer to the IRF in Equation (10) as $\bm{\delta}_{\tau,h}^{(d)}$ instead of $\bm{\delta}_{\tau,h}^{(d,0)}$.
Conventional IRFs from conditionally linear VARs are symmetric and proportional with respect to the sign and size of shocks, and they are time-invariant, at least in the case of constant parameter VARs. By contrast, in nonlinear models such as those we consider in this paper, the sign, size, and timing of a shock may matter and can lead to nonlinearities in responses. Due to this state dependence and the recursive nature of our model, we cannot use standard methods but need to resort to a variant of GIRFs instead, see Koop et al. (1996). There are multiple ways how one can define nonlinear IRFs, and we discuss our implementation of GIRFs below. We then extend the baseline framework to allow for studying various types of counterfactual shock and transmission scenarios.
Suppose we have (either internally or externally) identified the impact of a structural shock $u_{j\tau}$ of size $d$ on the endogenous variables (standard approaches from the SVAR literature, e.g., zero/sign restrictions, instrumental/proxy variables, can be used to identify the contemporaneous impact of the structural shocks). With a slight abuse of notation, define the impact as $(\bm{y}_{\tau}\>|\>u_{j\tau}=d_{0}+d)-(\bm{y}_{\tau}\>|\>,u_{j\tau}=d_{0})=d\bm{\beta}_{0}^{(m)}$ at iteration $m$. For simplicity, we only consider the typical baseline case $d_{0}=0$ in what follows (i.e., our notation here again abstracts from nonlinear shock impacts). [^4] Extensions to measure nonlinear shock impacts, see Caravello and Martinez-Bruera (2024); Gonçalves et al. (2024) for related discussions, are comparatively straightforward when treating the shock as observed and exogenous. Estimating nonlinear impacts for unobserved shocks — e.g., $\bm{\epsilon}_{t}=\bm{G}((u_{1t},\ldots,u_{nt})^{\prime})$, assuming that the reduced form shocks are nonlinear or unknown functions of the structural ones — complicates identification and inference, but is feasible subject to certain assumptions. To sample from the posterior distribution of the GIRFs across horizons, we will use the conditionally Gaussian form of Equation (4). Our goal here is to compare factual and counterfactual scenarios, i.e., we want to compare the case when a structural shock occurs with a non-shock baseline forecast. We construct these scenarios by first defining two distinct versions of initial conditions. In particular, let superscript $d$ indicate the shock scenario for each iteration $m$ of our algorithm, and set $\bm{x}_{\tau+1}^{(m,d)}=((\bm{y}_{\tau}+d\bm{\beta}_{0}^{(m)})^{\prime},\bm{y}_{\tau-1}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$, and $\bm{x}_{\tau+1}^{(m,\varnothing)}=(\bm{y}_{\tau}^{\prime},\bm{y}_{\tau-1}^{\prime},\ldots,\bm{y}_{\tau-p+1}^{\prime})^{\prime}$ with superscript $\varnothing$ referring to the no-shock scenario, i.e., the actual configuration of the input vector at time $\tau$ which yields the baseline prediction. This is similar in spirit to Gallant et al. (1993) and in line with how Gonçalves et al. (2021) define conditional IRFs.
Let $\bm{J}$ be a $k\times n$ selection matrix which has an identity matrix in its upper $n\times n$ block and zeroes everywhere else. The GIRF on impact reflects our desired shock scenario and is given by:
$\bm{\delta}_{\tau,0}^{(m,d)}=\bm{J}^{\prime}(\bm{x}_{\tau+1}^{(m,d)}-\bm{x}_{\tau+1}^{(m,\varnothing)})=d\bm{\beta}_{0}^{(m)}.$
That is, we assume that the impact of the respective shock is constant over time, and shocks of different signs and sizes are (proportionally) introduced by setting $d$ accordingly. Using Equation (4), we obtain two predictive distributions of interest:
$p(\bm{y}_{\tau+1}\>|\>u_{j\tau}=d,\mathcal{I},\bm{\Xi}^{(m)})$ $=\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,d)})+\bm{\epsilon}_{\tau+1},$
$p(\bm{y}_{\tau+1}\>|\>u_{j\tau}=0,\mathcal{I},\bm{\Xi}^{(m)})$ $=\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,\varnothing)})+\bm{\epsilon}_{\tau+1},$
where $\bm{\epsilon}_{\tau+1}\sim\mathcal{N}(\bm{0}_{n},\bm{\Sigma}_{\tau+1}^{(m)})$. In line with our IRF definition in Equation (10), at horizon $h=1$, we have $\bm{\delta}_{\tau,1}^{(m,d)}=\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,d)})-\bm{F}^{(m)}(\bm{x}_{\tau+1}^{(m,\varnothing)})$. [^5] Note that via our Monte Carlo approach we may also compute any potential functions of interest rather than the expectation, i.e., redefine Equation (10) as $g(\bm{y}_{\tau+h}\>|\>\mathcal{I},u_{j\tau}=d_{0}+d)-g(\bm{y}_{\tau+h}\>|\>\mathcal{I},u_{j\tau}=d_{0})$. The function $g(\bullet)$ could refer to, e.g., probabilities or quantiles, see also Gallant et al. (1993), and more recently, Jordà and Taylor (2024). For higher-order responses $h\geq 2$, we may iterate forward, sampling from the distribution of the future (reduced form) shocks to obtain draws from the unconditional predictive distribution and the distribution conditioning on a shock at time $\tau$ for each horizon. Specifically, for $s\in\{d,\varnothing\}$ and analogous to Equation (5), we have $\bm{x}_{\tau+h}^{(m,s)}=(\bm{y}_{\tau+h-1}^{(m,s)}{{}^{\prime}},\bm{y}_{\tau+h-2}^{(m,s)}{{}^{\prime}},\ldots)^{\prime}$ and thus,
$p(\bm{y}_{\tau+h}\>|\>u_{j\tau}=s,\mathcal{I},\{\bm{y}_{t}^{(m,s)}\}_{t=\tau+1}^{\tau+h-1},\bm{\Xi}^{(m)})=\mathcal{N}(\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,s)}),\bm{\Sigma}_{\tau+h}^{(m)}),$ (11)
which we use to compute the GIRF at generic horizon $h$:
$\bm{\delta}_{\tau,h}^{(m,d)}$ $=\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,d)})-\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,\varnothing)}).$ (12)
When $\bm{F}(\bullet)$ is linear, our approach produces “traditional” IRFs. This differentiates our version of GIRFs from the “original” GIRF. The latter also does not generally orthogonalize the shocks via SVAR approaches, see Koop et al. (1996); Pesaran and Shin (1998). By contrast, when $\bm{F}(\bullet)$ is nonlinear, the presence of these nonlinearities complicates the moments of the required higher-order responses, as discussed in Section 2. Revisiting the issue — for instance, taking $h=2$ as an example — when $\bm{F}$ is a linear mapping with associated (companion form) coefficients, $\tilde{\bm{A}}$, we obtain $\mathbb{E}(\tilde{\bm{A}}^{2}\bm{x}_{\tau+1}+\tilde{\bm{A}}\bm{\epsilon}_{\tau+1}+\bm{\epsilon}_{\tau+2})=\tilde{\bm{A}}^{2}\bm{x}_{\tau+1}$. That is, the IRF at horizon $h$ is given by $\bm{J}^{\prime}\tilde{\bm{A}}^{h}(\bm{x}_{\tau+1}^{(m,d)}-\bm{x}_{\tau+1}^{(m,\varnothing)})$ where $\bm{x}_{\tau+1}^{(m,d)}-\bm{x}_{\tau+1}^{(m,\varnothing)}=(d\bm{\beta}_{0}^{(m)}{{}^{\prime}},\bm{0}_{n(p-1)}^{\prime})^{\prime}$, and can be obtained by simply projecting the shock impact forward using powers of $\tilde{\bm{A}}$. Since $d$ can be factored out and the initial conditions cancel, the IRFs are time-invariant and proportional for different shock sizes and signs. In the more general nonlinear context, this is clearly not the case.
Equation (12) thus addresses the reduced form shocks that enter nonlinearly to obtain the respective conditional expectation (implicitly via the recursive simulation of the factual and counterfactual paths of the variables $\bm{y}_{\tau+1:\tau+h-1}^{(m,s)}$). Computing $\bm{\delta}_{\tau,h}^{(m,d)}$ for each iteration $m$ allows us to explore the posterior of the expression in Equation (10). Notably, we obtain a GIRF across horizons $h=0,1,2,\ldots,$ for each point in time $\tau$. In principle, one may thus consider “time-varying” dynamic effects of shocks for each period individually, or averages for different known subsets of observations (e.g., recessions vs. expansions). This time variation is due to the variation across initial conditions, and reflects the conditional IRF definition of Gonçalves et al. (2021). Unconditional IRFs, following the wording of this paper, can be computed in various ways (e.g., by randomizing over initial conditions and averaging, or by plugging in the unconditional moments of the variables as initial conditions). Our preferred approach, which we also use in our empirical applications, is to compute the conditional versions for each $\tau$. We subsequently take a time average, $\overline{\bm{\delta}}_{h}^{(m,d)}=1/T\sum_{\tau=1}^{T}\bm{\delta}_{\tau,h}^{(m,d)}$, to integrate out this source of randomness.
Another possibility that we explore in our applied work combines our approach to CFs with our algorithm for computing GIRFs. Specifically, we may switch off specific transmission channels of structural shocks by partially matching the moments of the predictive distribution conditional on the shock with those of the unconditional predictive distribution. By construction, this results in a GIRF that is equal to zero for the restricted dimensions, thereby effectively considering an alternative scenario where the shock cannot propagate through this channel. From an implementation perspective, we may manipulate Equation (11) for $s=d$ using a hard restriction as in Equation (9), such that $\bm{r}_{h}^{(m)}=\bm{R}_{h}\bm{F}^{(m)}(\bm{x}_{\tau+h}^{(m,\varnothing)})$. That is, we impose that $\bm{R}_{h}\bm{\delta}_{\tau,h}^{(m,d)}=\bm{0}_{r_{h}}$ along the desired $r_{h}$ dimensions.
3. Model Specification and Estimation Algorithm
3.1. Multivariate System Estimation
To estimate the multivariate (reduced form) model in Equation (1) efficiently from a computational perspective, we rely on a conditional representation of its $n$ equations. Let $\bm{e}_{i}$ of size $1\times n$ denote the $i$th row of an identity matrix $\bm{I}_{n}$, and $\bm{E}_{i}$ of size $(n-1)\times n$ is the matrix resulting from deleting the $i$th row of $\bm{I}_{n}$. Using $\bm{y}_{-it}=\bm{E}_{i}\bm{y}_{t}$ we may write $\bm{y}_{t}=\bm{e}_{i}^{\prime}y_{it}+\bm{E}_{i}^{\prime}\bm{y}_{-it}$. Under the assumptions of Equation (1) one may derive the conditional distribution: [^6] Related papers often either use a mapping between the structural and reduced form of the VAR to enable equation-by-equation estimation (see, e.g., Hauzenberger et al., 2024c), or rely on factor models for the reduced form errors (see, e.g., Clark et al., 2023). These approaches come with computational and inferential advantages and disadvantages. The former is simple to implement but requires parameterizing a structural form, which may cause issues such as inadvertently (instead of purposefully to achieve structural identification) breaking order-invariance of the equations. The latter allows for order-invariant inference but gives rise to the usual identification challenges of factor models. Our proposed approach uses a reduced form model, and we obtain order-invariant inference. See also Arias et al. (2023); Chan et al. (2024) for related discussions.
$p(y_{it}|\bm{y}_{-it},\bullet)\propto\exp\left\{-\frac{1}{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime}y_{it}^{2}-2y_{it}\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))\right\},$
which is a Gaussian with variance $\varsigma_{it}^{2}=(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{e}_{i}^{\prime})^{-1}$ and mean ${\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}(\bm{F}(\bm{x}_{t})-\bm{E}_{i}^{\prime}\bm{y}_{-it}))$. Indeed, this distribution is equivalent to a more common representation of the conditional multivariate Gaussian (see, e.g., Cong et al., 2017, Section 2). The mean can alternatively be written as ${\mu}_{it}=f_{i}(\bm{x}_{t})+\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, and we define $\tilde{\mu}_{it}=\varsigma_{it}^{2}(\bm{e}_{i}\bm{\Sigma}_{t}^{-1}\bm{E}_{i}^{\prime})(\bm{y}_{-it}-\bm{E}_{i}\bm{F}(\bm{x}_{t}))$, i.e., $\mu_{it}=f_{i}(\bm{x}_{t})+\tilde{\mu}_{it}$.
The $i$th equation of the multivariate model in regression form, conditional on all other equations, is then given by:
$(y_{it}-\tilde{\mu}_{it})=f_{i}(\bm{x}_{t})+u_{it},\quad u_{it}\sim\mathcal{N}(0,\varsigma_{it}^{2}),$ (13)
which can be used in a Gibbs sampler to update the conditional mean relationships by looping through equations $i=1,\ldots,n$. This approach is similar to the one of Esser et al. (2024) and allows to treat each equation of the multivariate system individually, conditional on all other equations. Specifics about how $f_{i}(\bm{x}_{t})$ is estimated, which we turn to next, are virtually irrelevant for using this equation-by-equation estimation algorithm.
3.2. Bayesian Additive Regression Trees
The approach we discuss in Section 2 works with any implementation of multivariate models with jointly Gaussian errors. That is, assuming a linear functional form for $\bm{F}(\bm{x}_{t})$ combined with suitable priors results in a standard BVAR. [^7] When we consider linear versions of our model for comparisons, we implement this setting with $\bm{F}(\bm{x}_{t})=\bm{A}\bm{x}_{t}$ where $\bm{A}$ is an $n\times k$ matrix of reduced form VAR coefficients. We assume a horseshoe prior with a single global shrinkage parameter on these coefficients, see also Hauzenberger et al. (2024b). Equation (13) can be used to update the VAR coefficients equation-by-equation from their usual Gaussian posteriors. In case we treat $\bm{F}(\bm{x}_{t})$ nonparametrically, several options are available. Due to its versatility and the established favorable empirical properties we mentioned earlier, we use BART to approximate the equation-specific functions in our applied work. That is, we consider a sum of $s=1,\ldots,S,$ tree functions $\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is})$:
$f_{i}(\bm{x}_{t})\approx\sum_{s=1}^{S}\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is}),$
where $\mathcal{T}_{is}$ are regression trees and $\bm{\mathrm{m}}_{is}$ is a vector of terminal node parameters (which serve as fitted values). Instead of having a single but complex tree, BART is akin to ensemble methods, and uses a sum of many simple trees (“weak learners”), which has been shown to work well empirically.
Using BART requires an algorithm that estimates splitting variables and thresholds for which we specify suitable priors that together yield $p(\mathcal{T}_{is})$; we further need a prior on the terminal node parameters $p(\bm{\mathrm{m}}_{is}|\mathcal{T}_{is})$. Our setup follows Chipman et al. (2010) and we first define the probability that a tree ends at a specific node at depth $d=0,1,2,\ldots,$ as $\alpha/(1+d)^{\beta}$, with $\alpha\in(0,1)$ and $\beta\in\mathbb{R}^{+}$. This prevents trees from getting overly complex and provides regularization (here, we rely on the default values $\alpha=0.95$, and $\beta=2$, which perform well across many datasets). For the splitting variables, we choose a uniform prior. This implies that each predictor is equally likely to be selected as a splitting variable. We further assign a uniform prior to all thresholds within the splitting rules, based on the range of the respective splitting variable.
Next we specify the prior for the terminal node parameters. On these parameters $\bm{\mathrm{m}}_{is,l}$, for $l=1,\ldots,\#\text{TN}_{is}$, where #$\#\text{TN}_{is}$ denotes the number of terminal node parameters of tree $s$ in equation $i$, we impose independent conjugate Gaussian priors that are symmetric across trees and identical for all terminal nodes. As suggested by Chipman et al. (2010) the moments of these priors are chosen in a data-driven manner, such that $95$% of the prior probability lies in the interval $(\min(\bm{y}_{i}),\max(\bm{y}_{i}))$, where $\bm{y}_{i}=(y_{i1},\ldots,y_{iT})^{\prime}$, and such that shrinkage increases the more trees $S$ are chosen for estimation. We choose $S=250$ trees which has been shown to work well for typical macroeconomic time series applications (see, e.g., Huber et al., 2023).
3.3. Priors on Other Model Parameters
The methods discussed in this paper work with a general time-varying covariance matrix $\bm{\Sigma}_{t}$. Significant computational advantages, however, are available if one assumes that $\bm{\Sigma}_{t}=s_{t}\bm{\Sigma}$, i.e., that the covariance structure only varies proportionally over time. The prior setup for the constant part of the covariance matrix follows Esser et al. (2024). Specifically, we use a hierarchical inverse Wishart prior:
$\bm{\Sigma}|\{a_{i}\}_{i=1}^{n}\sim\mathcal{W}^{-1}(s_{0},\bm{S}_{0}),$
where $s_{0}=\nu+n-1$, $\bm{S}_{0}=2\nu\cdot\text{diag}(1/a_{1},\ldots,1/a_{n})$ and $a_{i}\sim\mathcal{G}^{-1}(1/2,1/A_{j}^{2})$ for $i=1,\ldots,n,$ and a fixed scale parameter $A_{j}>0$; see also Huang and Wand (2013). Setting $\nu=2$ implies a comparatively uninformative prior about the implied correlation structure, different from fixed-hyperparameter versions of this prior.
In case we model time-varying variances, we follow Carriero et al. (2024) and assume that:
$s_{t}^{1/2}=\begin{cases}1&\text{with probability}\quad 1-\mathfrak{p}\\ \mathcal{U}(2,\overline{\mathfrak{s}})&\text{with probability}\quad\mathfrak{p},\end{cases}$
where $\mathcal{U}(2,\overline{\mathfrak{s}})$ is a discrete uniform distribution with (integer) support between $2$ and $\overline{\mathfrak{s}}=6$ and $\mathfrak{p}\sim\mathcal{B}(a_{\mathfrak{p}},b_{\mathfrak{p}})$ is the probability associated with observing an outlier. Alternative models related to common stochastic volatility specifications are available in this context. More flexible approaches, such as those discussed in Chan (2020, 2023), are straightforward to implement, but may significantly increase the computational burden.
3.4. Posterior Distributions and Sampling Algorithm
We may use the conditional distribution in Equation (13) to update the trees equation-by-equation using the backfitting approach designed by Chipman et al. (2010); see also Esser et al. (2024). Here, one may define the vector of partial residuals
$\tilde{y}_{is,t}=\left(y_{it}-\tilde{\mu}_{it}-\sum_{j\neq s}\ell_{ij}(\bm{x}_{t}\>|\>\mathcal{T}_{ij},\bm{\mathrm{m}}_{ij})\right)\sim\mathcal{N}\left(\ell_{is}(\bm{x}_{t}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is}),\varsigma_{it}^{2}\right),$
conditioning on the fit of each of the $S-1$ trees except tree $s$ and information in all but the $i$th equation. In full data notation, $\tilde{\bm{y}}_{is}=(\tilde{y}_{is,1},\ldots,\tilde{y}_{is,T})^{\prime}$, this defines a conditionally Gaussian likelihood, $p(\tilde{\bm{y}}_{is}\>|\>\mathcal{T}_{is},\bm{\mathrm{m}}_{is},\bullet)$, which can be marginalized analytically over the terminal node parameters $\bm{\mathrm{m}}_{is}$ (to keep the dimensionality of the inferential problem fixed).
Combining this conditional likelihood with the prior on the trees, and a suitable transition density (based on four distinct moves: grow a terminal node, prune a terminal node, change a splitting rule, swap a child/parent node), the trees are sampled using a standard accept/reject Metropolis-Hastings algorithm. These trees (and associated rules) partition the input space and we obtain a distinct set of observations for each terminal node. The posterior then takes the conventional Gaussian form for these parameters.
Updating all trees $s=1,\ldots,S,$ across equations $i=1,\ldots,n,$ yields an updated fit that can be used to compute the outlier-adjusted residuals $\bm{\epsilon}_{t}/\sqrt{s_{t}}=\bm{y}_{t}-\bm{F}(\bm{x}_{t})$. The posterior of the constant part of the covariance matrix is then given by:
$\bm{\Sigma}\>|\>\bullet\sim\mathcal{W}^{-1}\left(s_{0}+T,\bm{S}_{0}+\sum_{t=1}^{T}s_{t}^{-1}\bm{\epsilon}_{t}\bm{\epsilon}_{t}^{\prime}\right)$
The hierarchical parameters of the prior on the covariance matrix can be updated using:
$a_{i}\>|\>\bullet\sim\mathcal{G}^{-1}\left(\frac{\nu+T}{2},\frac{1}{A_{i}^{2}}+\nu\cdot\bm{\Sigma}^{-1}_{[ii]}\right),$
where $\bm{\Sigma}^{-1}_{[ii]}$ denotes the $i$th diagonal element of $\bm{\Sigma}^{-1}$. The outlier adjustment parameter $s_{t}$ can be sampled, due to its discrete support, using the probabilities:
$\Pr(s_{t}^{1/2}=1\>|\>\bm{y}_{t},\bm{F}(\bm{x}_{t}),\bm{\Sigma},\mathfrak{p})$ $\propto\mathcal{N}(\bm{y}_{t}\>|\>\bm{F}(\bm{x}_{t}),\bm{\Sigma})\cdot(1-\mathfrak{p}),$
$\Pr(s_{t}^{1/2}=\mathfrak{s}\>|\>\bm{y}_{t},\bm{F}(\bm{x}_{t}),\bm{\Sigma},\mathfrak{p})$ $\propto\mathcal{N}(\bm{y}_{t}\>|\>\bm{F}(\bm{x}_{t}),\mathfrak{s}\bm{\Sigma})\cdot\left(\mathfrak{p}/(\overline{\mathfrak{s}}-1)\right),\quad\text{for }\mathfrak{s}=2,3,\ldots,\overline{\mathfrak{s}},$
on a $t$-by-$t$ basis. The posterior distribution of the outlier probability is $\mathfrak{p}\>|\>\bullet\sim\mathcal{B}(a_{\mathfrak{p}}+T_{\text{o}},b_{\mathfrak{p}}+T-T_{\text{o}})$, with the total number of observations classified as outliers denoted by $T_{\text{o}}=\sum_{t=1}^{T}\mathbb{I}(s_{t}\neq 1)$, where $\mathbb{I}(\bullet)$ is an indicator function that yields $1$ if its argument is true and 00 otherwise. This completes our modeling framework and algorithmic implementation.
4. Empirical Applications
We employ the proposed framework in three related yet distinct applications. First, we use the annual stress test scenarios conducted by the Federal Reserve System and compute CFs using soft constraints on several variables. Second, we study the implications of varying financial conditions (imposing a hard constraint on a period) on tail risks of output growth, inflation, and employment. Third, we identify a US-based financial shock and gauge the role of spillovers and spillbacks to several other major economies. Across these applications we mainly compare a homoskedastic (BART-hom, setting $s_{t}=1$ for all $t$) and heteroskedastic BART implementation (BART-het) with a heteroskedastic BVAR (BVAR-het), both of which feature the outlier specification. Due to our sample featuring the Covid-19 pandemic observations we disregard the homoskedastic BVAR version (see Lenza and Primiceri, 2022, for a discussion).
4.1. Stress Testing Scenarios for the US Economy
In our first application, we conduct a scenario analysis for the US economy inspired by the 2024 version of the Dodd-Frank Act (DFA) stress test assumptions. This annual stress test is conducted and published by the Board of Governors of the Federal Reserve System. Details about the underlying dataset are provided in Appendix A. The information set features about $25$ broad variables (capturing economic activity, the labor market, prices, housing and the financial sector). We estimate our models using data on a quarterly frequency ranging from 1976Q1 to 2023Q4, and subsequently consider a baseline and adverse scenario for the period from 2024Q1 to 2027Q1. These scenarios are imposed via soft constraints on the future path of the unemployment rate (UNRATE), CPI inflation (CPIAUCSL), and $10$-year government bond yields (GS10), similar to Chan et al. (2025). They are visualized in Figure 2.
Figure 2: Restrictions imposed on the future paths of the indicated variables.
Notes: Scenarios according to the 2024 DFA stress test assumptions (solid lines), shaded areas indicate the bounds of the imposed soft restrictions. Variables: Consumer price inflation (CPIAUCSL), unemployment rate (UNRATE), and $10$-year government bond yields (GS10).
We display the posterior median forecasts alongside $50$ and $68$ percent posterior credible sets for selected unrestricted variables in Figure 3: real GDP (GDPC1), industrial production (INDPRO), personal consumption expenditure (PCE) inflation (PCECTPI), payroll employment (PAYEMS), $1$-year government bond yields (GS1) and the Gilchrist and Zakrajšek (2012) excess bond premium (EBP). The baseline scenario draws from the consensus projections from 2024 Blue Chip Financial Forecasts and Blue Chip Economic Indicators. The adverse scenario is characterized by a severe recession. It is noteworthy that the unconditional forecasts (in grey shades) approximately coincide with the baseline scenario (blue shades) in most cases. This is unsurprising, given that this scenario is designed to reflect the expectations of market participants, who apparently predict the economy to evolve similarly to what the models we consider here predict. Differences emerge irrespective of the linear vs. nonlinear specification for short-term interest rates, which are predicted to stay at a somewhat higher level unconditionally; and the BVAR suggests inflation to be elevated when comparing the unconditional forecast to the baseline scenario.
Turning to the adverse scenario, a different picture emerges — specifically, the scenario forecasts differ significantly from the unconditional forecasts. And there are some differences arising from nonlinearities in both the conditional means and variances. Interestingly, the different model specifications mostly agree about path of PCE inflation, financial conditions and the short-term interest rate conditional on the adverse scenario. As expected, financial conditions tighten initially but tend to improve subsequently, partially through a monetary easing response by the central bank as reflected in short-term interest rates. In addition, the assumed trajectory of the conditioning variables results in a modestly disinflationary episode that vanishes by $2026$.
(a) BART-hom
(b) BART-het
(c) BVAR-het
Figure 3: Conditional and unconditional forecasts for selected variables.
Notes: Posterior median alongside 50/68 percent credible sets. Variables: Real GDP (GDPC1), industrial production (INDPRO), personal consumption expenditure inflation (PCECTPI), payroll employment (PAYEMS), $1$-year government bond yields (GS1) and excess bond premium (EBP).
Major differences across model implementations emerge when considering measures of economic activity and payroll employment. While all three models agree about the fact that the adverse scenario leads to a significant economic contraction, the heteroskedastic BART model suggests somewhat less severe magnitudes of this downturn. This can be explained by noting that for this specification, the algorithm decides to classify several observations (that are otherwise informative about directional movements of variables when assuming homoskedasticity) as more noisy, or even as outliers. We note that BART, due to the way how tree-based approaches fit data, is capable of dealing with outliers and heteroskedastic data features in the conditional mean function by design (see Huber et al., 2023; Clark et al., 2023, for discussions), even when assuming homoskedastic reduced form errors. But in this case it classifies several observations instead as noise rather than signal.
Comparing the homoskedastic BART-version with the BVAR, which are closer in terms of conditional predictive distributions for the adverse scenario, both suggest a significant decline in industrial production, with a trough at $-10$ percent (about $-6$ percent for heteroskedastic BART). Relatedly, the BVAR predicts a decline in real GDP of about $-4$ percent, roughly in line with the numbers obtained when running homoskedastic BART, at $-5$ percent. Where the two differ in this context is in the predicted recovery following the trough, which is more sluggish for the nonparametric model. Turning to payroll employment, homoskedastic BART suggests a sharper decline of about $-6$ percent, which is about twice the size of the decline conditionally forecasted by the heteroskedastic BART and BVAR models.
4.2. Financial Conditions in the US and Tail Risk Scenarios
In this empirical application, we restrict our sample from 1976Q1 to 2017Q4 and consider the period from 2018Q1 until 2019Q1 as a laboratory to assess nonlinearities in the relationship between economic variables and financial conditions. For this application we use the homoskedastic BART model as the pandemic observations are excluded from our dataset. We investigate nonlinear patterns of macroeconomic risk, which, following the “growth-at-risk” approach of Adrian et al. (2019), is defined as the predictive quantiles of some variable of interest at a pre-defined probability level (in line with value-at-risk, VaR, in finance). We pick this period because the information set already contains the global financial crisis (and the model thus had the opportunity to learn from this severe financial episode), and because this “holdout sample” otherwise coincides with a comparatively eventless period. We impose hard constraints on the future path of the National Financial Conditions Index (NFCI) and trace the effects of these scenarios on several macroeconomic variables.
The scenarios are defined to reflect an increase of the NFCI ranging from approximately $1$, $3$ and $6$ unconditional standard deviations (reflecting tighter financial conditions) in 2018Q1, i.e., in $\mathcal{C}_{1}$, which we implement by placing these values as hard restriction on the NFCI in that quarter. From 2018Q2 onward we leave the respective future path unrestricted, i.e., $\mathcal{C}_{h}=\varnothing$ for $h>1$. It is worth mentioning that in this specific case the two possible variants of CF as discussed in Section 2.2 are identical.
We investigate growth-at-risk (quantiles of real GDP), inflation-at-risk (quantiles of PCE inflation) and labor-at-risk (quantiles of growth in payroll employment) as our objects of interest (see Adams et al., 2021; Pfarrhofer, 2022; Botelho et al., 2024; Clark et al., 2024; Lopez-Salido and Loria, 2024, for related papers). The resulting CF distributions (density estimates) are shown in Figure 4. Computing the difference between these conditional scenario distributions and the unconditional one yields an estimate of the unorthogonalized IRF (UIRF). Reflecting Equation (10), we have $\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I},\mathcal{C}_{1})-\mathbb{E}(\bm{y}_{\tau+h}\>|\>\mathcal{I})$, which is the counterfactual difference between the scenario defined by the restriction in $\mathcal{C}_{1}$ and the unconditional forecast. Different to a simulation of an individual structural shock, the UIRF reflects a likely combination of structural shocks that increases the NFCI by the predefined amount, see also Crump et al. (2025).
Figure 4: Conditional forecast distributions for selected variables and macroeconomic value-at-risk (VaR) for different scenarios of financial stress.
Notes: The crosses mark the posterior median for unconditional forecasts and circles indicate realized values. Max. NFCI refers to the maximum value of the NFCI for the scenarios, with moderate (1), severe (3, comparable to the global financial crisis), and extreme stress (6). Sampling period 1976Q1 to 2017Q4, the hard restriction applies for 2018Q1. Unorthogonalized impulse response functions (UIRFs) are computed as the difference between the conditional and unconditional distribution (posterior median alongside 50 and 68 percent credible sets). Variables: Real GDP (GDPC1), personal consumption expenditure inflation (PCECTPI), payroll employment (PAYEMS), national financial conditions index (NFCI).
The different NFCI scenarios for 2018Q1 shift the predictive distributions, especially during the quarter when this reduced form shock materializes. The resulting dynamic effects, as measured with the UIRFs, are rather short-lived for real GDP and PCE inflation, but more persistent for payroll employment. In all shown cases, the economy contracts which is reflected in a decrease of real GDP growth and payroll employment, and the simulated shock has a deflationary effect. A clear pattern that emerges is that while the upper tails of the distributions remain rather stable (upside risk is constant), downside risk as measured by the lower quantiles increases significantly for all considered variables (the red shaded VaR $<0.05$ moves strongly leftwards), and there are some visible asymmetries. While the moderate NFCI scenario (max. NFCI $=1$) results in growth-at-risk for the $5$th percentile at about $-5$ percent, the severe (max. NFCI $=3$) and extreme (max. NFCI $=6$) stress scenarios yield $-10$ and $-20$ percent, respectively. This finding is also present for inflation-at-risk and labor-at-risk.
The resulting predictive distributions exhibit non-Gaussian features, chief among them being heavy tails and skewness (see also Clark et al., 2023, for a related but simplified scenario analysis in this context). In addition, there are hints of multimodality as the assumed values for the NFCI in 2018Q1 turn more extreme, which relates to the discussions in Adrian et al. (2021). For instance, while most mass of the predictive distribution of payroll employment growth, for the extreme stress scenario, is centered on about $-3$ percent, there is a smaller local peak at values of $-5$ percent. A similar pattern is visible for PCE inflation. These features can arise — even in one-step-ahead predictions and for a single restricted period — due, for example, to the initial conditions of the nonlinear unconditional mean function at the time the forecast is generated.
4.3. Spillovers and Spillbacks of US Financial Shocks
This application estimates the effect of a financial shock in the US and traces its effects through the domestic economy, but also captures spillovers and spillbacks to and from several other economies. We use a slightly different dataset in this case, which drops several of the domestic indicators, but adds bilateral exchange rates alongside real GDP for the EA, Japan and the UK. Following the linear SVAR literature, we assume that the reduced form shocks are linked to the structural shocks as $\bm{\epsilon}_{t}=\bm{B}_{0}^{-1}\bm{u}_{t}$. This implies that $\bm{\Sigma}=\bm{B}_{0}^{-1}\bm{B}_{0}^{-1}{{}^{\prime}}$ and that shocks impact the dynamic model linearly. Suppose the structural shock of interest is in the $j$th position of $\bm{u}_{t}$ and $\bm{u}_{-jt}=\bm{0}_{n-1}$, then the impact response is $d\bm{B}_{0}^{-1}\bm{e}_{j}^{\prime}=d\bm{\beta}_{0}$. That is, $\bm{\beta}_{0}$ is the $j$th column of the matrix $\bm{B}_{0}^{-1}$.
We identify $\bm{B}_{0}^{-1}$ using a set of zero impact restrictions imposed with a Cholesky decomposition of the form $\bm{\Sigma}=\bm{D}\bm{D}^{\prime}$ where $\bm{D}$ is lower triangular. That is, $\bm{B}_{0}^{-1}=\bm{D}$, and we orthogonalize the structural shocks (different to the application in Section 4.2 which dealt with unorthogonalized shocks that resulted in an increase of the NFCI). We simulate different shock sizes of different signs with $d\in\{-3,-1,1,3,6\}$. Different to conventional linear frameworks, our approach allows to assess nonlinearities of higher-order responses with respect to different signs and magnitudes of a proportional shock impact with GIRFs. Such asymmetries and related nonlinearities have recently gained attention both in a VAR and local projection context, see, e.g., Mumtaz and Piffer (2022); Carriero et al. (2023); Forni et al. (2024); Hauzenberger et al. (2024a).
To identify the financial shock, we place timing-restrictions on the contemporaneous impulse responses as described above. This is operationalized with a specific ordering of the quantities in the vector $\bm{y}_{t}$ — we structure this vector such that all slow moving domestic and foreign macroeconomic variables come first (which imposes zero restrictions on impact). These variables are then followed by the excess bond premium (EBP, Gilchrist and Zakrajšek, 2012), and all fast moving variables such as those capturing the financial economy. We interpret the orthogonalized innovation of the EBP equation as the financial shock, similar to Barnichon et al. (2022). In a multicountry context, Huber et al. (2024) use an identification scheme virtually identical to ours.
Figure 5: Impulse response functions for selected variables to a financial shock in the US, comparing asymmetries due to size and sign of the shocks.
Notes: Posterior median alongside 50 and 68 percent credible sets. Cumulated responses for variables in differences and levels for all other variables. Variables: Real GDP (GDPC1), consumer price inflation (CPIAUCSL), payroll employment (PAYEMS), federal funds rate (FEDFUNDS) and S&P500 index (SP500).
We present selected results for key macroeconomic and financial variables in Figure 5, additional empirical results can be found in the Appendix. While our framework allows to compute GIRFs for each period in our sample, we focus on time averages in the results that follow. The rows in the figure show different subsets of the same set of results, structured such that the shocks of different signs and sizes can be compared with ease. That is, for visualization purposes, we typically rescale the GIRFs to a common range by computing “normalized” GIRFs, $\tilde{\bm{\delta}}_{\tau,h}^{(m,d)}=\bm{\delta}_{\tau,h}^{(m,d)}/d$. Note that for linear VARs, such scaling yields identical IRFs for all shock sizes (see the additional results for the linear BVAR in the Appendix). For nonlinear models, this is not necessarily the case and allows for a visual inspection of asymmetries.
Starting with the first row in Figure 5, we find that the size of the financial shock does not matter much for the GIRFs about any of the indicated variables. Financial shocks of different sizes rescaled to reflect a $1$ standard deviation innovation rather symmetrically decrease real GDP and payroll employment. Peak effects occur between one and two years after impact of the shock. In addition, the shock puts a persistent downward pressure on prices and leads to a decline in the Federal funds rate which peaks at about $-10$ basis points after around a year. Notably it does not react on impact, different to stock returns which immediately decline by about $1.5$ percent during the quarter that the shock materializes. Qualitatively and in terms of magnitudes, these estimates are roughly in line with those in Gilchrist and Zakrajšek (2012). We do not report these results here to save space, but note that the US-based financial shock spills over to the other economies and leads to contractionary effects in terms of real GDP.
Having established that the size of the financial shock does not seem to matter much, the second and third row zoom into sign asymmetries. Here, we show the raw GIRF in grey, while a version of the normalized GIRF flips the sign to ease comparisons. The $1$ standard deviation US-based financial shock does not result in significant sign asymmetries (the posterior distributions overlap for the most part). Turning to the final row, this clearly differs for larger sized shocks of different signs. For these GIRFs that show the responses to a positive and negative $3$ standard deviation shock, asymmetries are striking and significant for most variables. In particular, we find that the negative effect on payroll employment is almost twice as large for adverse shocks, and the Federal Reserve responds more strongly to adverse financial shocks, as measure with the much stronger shift in the Federal funds rate. These findings corroborate the empirical evidence presented by the preceding literature (see, e.g., Forni et al., 2024; Hauzenberger et al., 2024a).
Figure 6: Unrestricted and restricted impulse response functions for selected variables.
Notes: The restricted case assumes that the US financial shock does not spill over to non-domestic variables. The upper panels show the GIRFs for these two cases, and the lower panel shows the posterior distribution of the differences between the unrestricted and no-spillovers scenario. Posterior medians alongside 50 percent credible sets. Cumulated responses for variables in differences and levels for all other variables.
We next explore the role of international variables in the domestic transmission of the US shock. For this purpose, besides unrestricted GIRFs, we consider an alternative analysis where non-domestic transmission channels are switched off. This is different to Huber et al. (2024), who focus on international effects of US-based shocks; indeed, we investigate how international channels affect the domestic transmission of shocks originating in the US. That is, we impose the restriction that foreign variables do not respond to the financial shock in the US in this counterfactual, thereby simulating a scenario where the financial shock is fully confined to the domestic economy without any spillovers (or spillbacks); see Breitenlechner et al. (2022) for a monetary application in this context.
The results are displayed in Figure 6. The upper panels show the unrestricted GIRFs (those shown and discussed in the context of Figure 5) and restricted “no spillovers” GIRFs. The lower panels plot the difference between the two. Two key findings are worth reporting. First, for the most part, ruling out spillovers and spillbacks does not significantly alter the GIRFs for most variables. This can be observed from the credible sets covering the zero line in all but one panel in the bottom row of the figure. Second, international transmission channels appear to matter for inflation dynamics after financial shocks. We find that in the no-spillover case, the restricted GIRFs to the shocks of different signs are clearly asymmetric. The inflation response is insignificant for benign financial shocks, but significantly negative for adverse shocks; this is different for unrestricted GIRFs, which are mostly symmetric.
In terms of the importance of international transmission channels for the domestic shock propagation, we find that non-domestic dynamics partially offset disinflationary or deflationary pressures in response to adverse US financial shocks. By contrast, in the no-spillover case, the inflation GIRFs take smaller or even negative values for benign financial shocks than in the unrestricted case. These patterns suggests that propagation through non-domestic variables overall lowers the inflation responses irrespective of the sign of the shock. In other words, ruling out international channels loosely speaking introduces a negative “bias” of the response of US inflation in this counterfactual simulation.
5. Conclusions
This paper presents a robust and unified methodology for conducting scenario analysis in multivariate macroeconomic settings, accommodating nonlinearities and unknown functional forms of conditional mean relationships. The proposed methods are applicable not only in traditional nonlinear frameworks, such as variants of threshold or time-varying parameter models, but also to more recently developed models incorporating machine learning techniques. By extending CFs and GIRFs to a flexible, nonparametric framework, we address some limitations of traditional linear and parametric models in generating various types of counterfactual analyses. Specifically, we explore the application of CFs in a nonlinear setting by leveraging the properties of conditionally Gaussian errors. Furthermore, we demonstrate how to derive dynamic causal effects in the form of GIRFs in a nonlinear context and quantify the contribution of specific transmission channels in the propagation of structural shocks. This approach is broadly applicable and computationally efficient, making it suitable for large-scale macroeconomic datasets.
Empirical applications, focusing on the use of BART as an example of nonparametric modeling of the conditional mean function, underscore the critical role of nonlinearities and heteroskedasticity in shaping macroeconomic dynamics. For instance, a scenario analysis based on Federal Reserve stress test assumptions reveals differences between linear and nonlinear models in forecasting economic contractions and recoveries. Similarly, our growth-at-risk application demonstrates how nonlinearities influence the distribution of macroeconomic risks under financial stress, particularly in amplifying downside risks. Finally, the analysis of financial spillovers reveals significant asymmetries in the transmission of shocks and the influence of international linkages on domestic outcomes. Our results indicate that incorporating nonlinearities in CFs and scenario analysis provides a richer understanding of risk propagation and policy effects. By addressing limitations in traditional linear approaches and offering a flexible tool for analyzing complex economic relationships, this paper contributes to the literature on macroeconomic forecasting, risk assessment, and policy evaluation.
Appendix A Data
Table A.1: Dataset. Notes: Variable codes, descriptions and transformation: (0) no transformation $h(x_{t})=x_{t}$; (1) annualized log-differences $h(x_{t})=400\cdot\log(x_{t}/x_{t-1})$; (2) log-differences $h(x_{t})=100\cdot\log(x_{t}/x_{t-1})$. Check marks indicate inclusion in the information sets for our applications (CF for Sections 4.1 and 4.2, GIRF for Section 4.3).
Code Description $h(x_{t})$ CF GIRF
GDPC1 Real gross domestic product 1
PCECC96 Real personal consumption expenditure 1
PRFIx Real private fixed investment (residential) 1
GCEC1 Real Government consumption and investment 1
RDI Real disposable income 0
INDPRO Industrial production 1
CPIAUCSL Headline consumer price index (CPI) 1
CPILFESL Core CPI (excl. food and energy) 1
PCECTPI Headline PCE prices 1
PCEPILFE Core PCE prices (excl. food and energy) 1
HPI House price index 1
HOUST Housing starts 1
MR Mortgage rate 0
PAYEMS Payroll employment 1
UNRATE Unemployment rate 0
FEDFUNDS Federal funds rate 0
GS1 1-year treasury rate 0
GS10 10-year treasury rate 0
EBP Excess bond premium 0
NFCI National financial conditions index 0
EXUSUKx US/UK foreign exchange rate 0
OILPRICEx Real crude oil prices (WTI) 2
SP500 S&P 500 2
EARGDP Real gross domestic product (EA) 0
JPRGDP Real gross domestic product (JP) 0
UKRGDP Real gross domestic product (UK) 0
USDEUR EU/US foreign exchange rate 0
JPUSD JP/US foreign exchange rate 0
Appendix B Additional Empirical Results
(a) BART-het
(b) BVAR-het
Figure B.1: Comparison of impulse response functions across nonlinear and linear models. Posterior medians alongside 68 percent credible sets. Cumulated responses for variables in differences and levels for all other variables.